OpenStudy (lesli):

what is the equation of the line that passes through (-2 3) and is parallel to 2x+3y=6

OpenStudy (triciaal):

parallel lines have the same slope put the given eqn in the form y=mx + b the slope is m the line parallel has this value for m and passes through the given point the parallel line will have 2 points (x,y) and the one given the point-slope eqn of the line is (y-y1) = m(x-x1) plug the numbers and simplify rewrite in standard form Ax + By = C